# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ II.1 Scenarios I and II
We now turn to the second scenario, scenario II, where the light anti-quarks are symmetric. For the $q\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}$ configuration, we have eighteen SU(3) ${\bf 3}\otimes\bar{\bf 6}$ states. They form a $\overline{\bf 15}$ and another $\bar{\bf 3}$, see Eq. (133). The $\overline{\bf 15}$ is the traceless part, which consists of six iso-multiplets: an $S=+1$ isotriplet, $(T^{++}_{c\bar{s}},T^{+}_{c\bar{s}},T^{0}_{c\bar{s}})$, an $S=-1$ isotriplet, $(T^{+}_{cs},T^{0}_{cs},T^{-}_{cs})$, an $S=+2$ isodoublet, $(T^{++}_{c\bar{s}\bar{s}},T^{+}_{c\bar{s}\bar{s}})$, an $S=0$ iso-quarplet, $(T_{c}^{++},T_{c}^{+},T_{c}^{0},T_{c}^{-})$, an $S=0$ isodoublet $(T_{c}^{\prime+},T_{c}^{\prime 0})$ and an $S=+1$ isosinglet, $T_{c\bar{s}}^{\prime+}$, while the $\bar{\bf 3}$ is the traceful part, and it has an $S=0$ isodoublet, $(T^{\prime\prime+}_{c},T^{\prime\prime 0}_{c})$, and an $S=+1$ isosinglet, $T^{\prime\prime+}_{c\bar{s}}$, see Fig. 1 (b). The quark contents and quantum numbers of these states are given in Table 2, and as in scenario I, we label these $T$ in Table 2 and add asterisks to flavor exotic states.
Note that in this scenario, $T^{*}_{c\bar{s}0}(2900)^{++}$, $T_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ are in the $\overline{\bf 15}$. In particular, $T^{*}_{c\bar{s}0}(2900)^{++}$ and $T_{c\bar{s}0}(2900)^{0}$ are $T^{++}_{c\bar{s}}$ and $T^{0}_{c\bar{s}}$, respectively, belonging to the $S=1$ isotriplet, while $T^{*}_{cs0}(2870)^{0}$ is $T^{0}_{cs}$ belonging to another isotriplet with $S=-1$. In this scenario, however, there are six other flavor exotic states similar to the above states, namely, $T^{+}_{cs}$. $T^{-}_{cs}$, $T^{++}_{c\bar{s}\bar{s}}$, $T^{+}_{c\bar{s}\bar{s}}$, $T^{++}_{c}$, and $T^{-}_{c}$. In fact, $T^{+}_{cs}$, $T^{-}_{cs}$ and $T^{*}_{cs0}(2870)^{0}$ belong to the same isotriplet (with $S=-1$), while $T^{++}_{c\bar{s}\bar{s}}$, $T^{+}_{c\bar{s}\bar{s}}$, $T^{++}_{c}$ and $T^{-}_{c}$ belong to the $S=-2$ isodouble and the iso-quarplet with $S=0$, respectively, as shown in Table 2. These nine flavor exotic states are labeled with asterisks in the table. They have been identified in the literature, see, for example, ref. Qin:2022nof , where a different notation was used.
Furthermore, as mentioned previously, mixing for particles in different isospin multiplets is expected to be more difficult to occur. Consequently, $T^{+}_{c\bar{s}}$ in the $S=+1$ isotripet, $T^{+}_{c}$ and $T^{0}_{c}$ in the $S=0$ iso-quarplet do not mix with $T^{\prime\prime+}_{c\bar{s}}$, $D_{s}$, $T^{\prime+}_{c}$, $T^{\prime 0}_{c}$ or $D^{+,0}$. They should also be considered flavor exotic particles and be labeled with asterisks in the table. To our knowledge, these particles have not been considered flavor exotic particles in the literature. Altogether, we have twelve flavor exotic states in this scenario.
@@
where only the $\bar{\bf 3}^{\prime}$ states remain after taking the trace, $T^{i}_{\{ik\}}$, as the $\overline{\bf 15}$ is traceless, where one can also see from Table 2.
@@ II.2 Remarks
In this work, we consider $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ either in the ${\bf 6}$ as in scenario I or in the $\overline{\bf 15}$ as in scenario II. In principle, they may belong to different multiplets. For example, we may have $T_{cs0}(2870)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{0}$ belonging to the ${\bf 6}$, while $T_{cs0}(2870)^{0}$ belonging to the $\overline{\bf 15}$, or vice verse. It is our working assumption that they only belong to one of the multiplets and work out the consequences. It is then useful to verify those predictions experimentally to check the working assumption. Therefore, we should pay attention to predictions where all $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ are involved, or, at least, $T_{cs0}(2870)^{0}$ and either one of $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ are involved.
Before we end this section, it should be pointed out that, in principle, particles in scenarios I and II can mix. It is our working assumption that they do not. Even if they do, the amplitudes obtained in this work are still useful, since amplitudes for mixed states can be built upon them. Nevertheless, not all flavor exotic states can mix. Although $T^{++}_{c\bar{s}}$, $T^{0}_{c\bar{s}}$ and $T^{0}_{c\bar{s}}$ in scenarios I and II can, $T_{cs}^{0}$s cannot, as they have different isospin quantum numbers. In addition, the rest of the flavor exotic states in scenario II cannot, as they do not have any counterparts in scenario I, see Fig. 1. Furthermore, all flavor exotic states identified in this section do not mix with $D$ or any other $c\bar{q}$ states in any case. They are exotic.
@@ III.1 $T\to DP$ decays
We consider the strong decays of $T\to DP$ decays in this sub-section. We will first discuss how to decompose these decay amplitudes into topological amplitudes. Then $T\to DP$ decay amplitudes will be decomposed accordingly.
@@ III.1.1 Topological amplitudes in $T\to DP$ decays
As shown in the above effective Hamiltonian, there are three topological amplitudes in $T\to DP$ decays, where $F_{DP}$ is a fall-apart amplitude, $AC_{P}$ and $AC_{D}$ are annihilation and creation amplitudes. The topological amplitudes are depicted in Fig. 2. The subscripts indicate the final state meson(s) receiving the light quark and/or the anti-quarks of $T_{cq\bar{q}^{\prime}\bar{q}^{\prime\prime}}$. For example, for $AC_{P}$, the antiquark of $T$ goes into $P$, see the diagram in Fig. 2(b). It is reasonable to expect that the fall-apart amplitude is dominating, i.e.
$|F_{DP}|\gg|AC_{P}|,|AC_{D}|.$ (25)
Furthermore, the decay amplitudes of tetraquarks in the ${\bf 6}$ in scenario I and in the $\overline{\bf 15}$ in scenario II can only have contributions from diagram (a), the fall-apart amplitude, $F_{DP}$. This can be easily seen, as the $AC_{P}$ and $AC_{D}$ terms in Eqs. (16) and (17) necessary involve the trace $T^{i}_{[ik]}$ or $T^{i}_{\{ik\}}$, where only the states in the $\bar{\bf 3}$ in scenario I and the $\bar{\bf 3}^{\prime}$ in scenario II remain, see Eqs.(11) and (15). In fact, the reason for the above facts can be easily understood by noting $\bar{\bf 3}\otimes{\bf 8}=\overline{\bf 15}\oplus{\bf 6}\oplus\bar{\bf 3}$, see Eq. (134). Hence there is only one $\overline{\bf 15}$ and one ${\bf 6}$ in $D\Pi$ final states, which can match the $\overline{\bf 15}$ of $T$ in scenario II and ${\bf 6}$ of $T$ in scenario I, respectively. Adding $\eta_{1}$ in $P$ does not affect the above matching.
@@ III.1.2 $T\to DP$ decay amplitudes
$1^{*}$ $T^{++}_{c\bar{s}}\to D^{+}K^{+}$ $F_{DP}$ $T^{++}_{c\bar{s}}\to D_{s}^{+}\pi^{+}$ $-F_{DP}$
$2^{*}$ $T^{+}_{c\bar{s}}\to D^{0}K^{+}$ $\frac{1}{\sqrt{2}}F_{DP}$ $T^{+}_{c\bar{s}}\to D^{+}K^{0}$ $-\frac{1}{\sqrt{2}}F_{DP}$
@@
$3^{*}$ $T^{0}_{c\bar{s}}\to D^{0}K^{0}$ $F_{DP}$ $T^{0}_{c\bar{s}}\to D_{s}^{+}\pi^{-}$ $-F_{DP}$
$4^{*}$ $T^{0}_{cs}\to D^{0}\overline{K}^{0}$ $F_{DP}$ $T^{0}_{cs}\to D^{+}K^{-}$ $-F_{DP}$
5 $T^{+}_{c}\to D^{0}\pi^{+}$ $\frac{1}{\sqrt{2}}F_{DP}$ $T^{+}_{c}\to D^{+}\pi^{0}$ $-\frac{1}{2}F_{DP}$
$T^{+}_{c}\to D_{s}^{+}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}F_{DP}$ $T^{+}_{cu\bar{u}\bar{d}}\to D^{+}\eta$ $-\frac{c_{\phi^{\prime}}+\sqrt{2}s_{\phi^{\prime}}}{2}F_{DP}$
$T^{+}_{c}\to D^{+}\eta^{\prime}$ $\frac{\sqrt{2}c_{\phi^{\prime}}-s_{\phi^{\prime}}}{2}F_{DP}$
6 $T^{0}_{c}\to D^{0}\pi^{0}$ $-\frac{1}{2}F_{DP}$ $T^{0}_{c}\to D^{+}\pi^{-}$ $-\frac{F_{DP}}{\sqrt{2}}$
$T^{0}_{c}\to D_{s}^{+}K^{-}$ $\frac{1}{\sqrt{2}}F_{DP}$ $T^{0}_{cd\bar{u}\bar{d}}\to D^{+}\eta$ $\frac{c_{\phi^{\prime}}+\sqrt{2}s_{\phi^{\prime}}}{2}F_{DP}$
$T^{0}_{c}\to D^{+}\eta^{\prime}$ $-\frac{\sqrt{2}c_{\phi^{\prime}}-s_{\phi^{\prime}}}{2}F_{DP}$
7 $T^{\prime\prime+}_{c}\to D^{0}\pi^{+}$ $\frac{1}{\sqrt{2}}(F_{DP}+2AC_{P})$ $T^{\prime\prime+}_{c}\to D^{+}\pi^{0}$ $-\frac{1}{2}(F_{DP}+2AC_{P})$
@@
$T^{\prime\prime+}_{c}\to D^{+}\eta$ $\frac{c_{\phi^{\prime}}+\sqrt{2}s_{\phi^{\prime}}}{6}(F_{DP}+2AC_{P})$ $T^{\prime\prime+}_{c}\to D^{+}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}-\sqrt{2}c_{\phi^{\prime}}}{6}(F_{DP}+2AC_{P})$
$+\frac{\sqrt{2}s_{\phi^{\prime}}-2c_{\phi^{\prime}}}{3}(F_{DP}-AC_{P}-3AC_{P})$ $-\frac{\sqrt{2}c_{\phi^{\prime}}+2s_{\phi^{\prime}}}{3}(F_{DP}-AC_{P}-3AC_{P})$
@@
$T^{\prime\prime 0}_{c}\to D^{0}\eta$ $\frac{c_{\phi^{\prime}}+\sqrt{2}s_{\phi^{\prime}}}{6}(F_{DP}+2AC_{P})$ $T^{\prime\prime 0}_{c}\to D^{+}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}-\sqrt{2}c_{\phi^{\prime}}}{6}(F_{DP}+2AC_{P})$
$+\frac{\sqrt{2}s_{\phi^{\prime}}-2c_{\phi^{\prime}}}{3}(F_{DP}-AC_{P}-3AC_{P})$ $-\frac{\sqrt{2}c_{\phi^{\prime}}+2s_{\phi^{\prime}}}{3}(F_{DP}-AC_{P}-3AC_{P})$
9 $T^{\prime\prime+}_{c\bar{s}}\to D^{0}K^{+}$ $\frac{1}{\sqrt{2}}(F_{DP}+2AC_{P})$ $T^{\prime\prime+}_{c\bar{s}}\to D^{0}K^{+}$ $\frac{1}{\sqrt{2}}(F_{DP}+2AC_{P})$
$T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\eta$ $-\frac{c_{\phi^{\prime}}+\sqrt{2}s_{\phi^{\prime}}}{3}(F_{DP}+2AC_{P})$ $T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\eta^{\prime}$ $\frac{\sqrt{2}c_{\phi^{\prime}}-s_{\phi^{\prime}}}{3}(F_{DP}+2AC_{P})$
$+\frac{\sqrt{2}s_{\phi^{\prime}}-2c_{\phi^{\prime}}}{3}(F_{DP}-AC_{P}-3AC_{P})$ $-\frac{\sqrt{2}c_{\phi^{\prime}}+2s_{\phi^{\prime}}}{3}(F_{DP}-AC_{P}-3AC_{P})$
@@
$2^{\prime*}$ $T^{+}_{c\bar{s}}\to D^{0}K^{+}$ $\frac{1}{\sqrt{2}}F_{DP}$ $T^{+}_{c\bar{s}}\to D^{+}K^{0}$ $-\frac{1}{\sqrt{2}}F_{DP}$
@@
$11^{\prime*}$ $T^{0}_{c}\to D^{0}\pi^{0}$ $\frac{2}{\sqrt{3}}F_{DP}$ $T^{0}_{c}\to D^{+}\pi^{-}$ $-\sqrt{\frac{2}{3}}F_{DP}$
@@
$13^{\prime}$ $T^{\prime+}_{c}\to D^{0}\pi^{+}$ $\frac{1}{2\sqrt{3}}F_{DP}$ $T^{\prime+}_{c}\to D^{+}\pi^{0}$ $-\frac{1}{2\sqrt{6}}F_{DP}$
$T^{\prime+}_{c}\to D^{+}\eta$ $\frac{\sqrt{6}}{4}(c_{\phi^{\prime}}+\sqrt{2}s_{\phi^{\prime}})F_{DP}$ $T^{\prime+}_{c}\to D^{+}\eta^{\prime}$ $\frac{\sqrt{6}}{4}(s_{\phi^{\prime}}-\sqrt{2}c_{\phi^{\prime}})F_{DP}$
$T^{\prime+}_{c}\to D_{s}^{+}\overline{K}^{0}$ $-\frac{\sqrt{3}}{2}F_{DP}$
$14^{\prime}$ $T^{\prime 0}_{c}\to D^{0}\pi^{0}$ $\frac{1}{2\sqrt{6}}F_{DP}$ $T^{\prime 0}_{c}\to D^{0}\eta$ $\frac{\sqrt{6}}{4}(c_{\phi^{\prime}}+\sqrt{2}s_{\phi^{\prime}})F_{DP}$
$T^{\prime 0}_{c}\to D^{0}\eta^{\prime}$ $\frac{\sqrt{6}}{4}(s_{\phi^{\prime}}-\sqrt{2}c_{\phi^{\prime}})F_{DP}$ $T^{\prime 0}_{c}\to D^{+}\pi^{-}$ $\frac{1}{2\sqrt{3}}F_{DP}$
$T^{\prime 0}_{c}\to D_{s}^{+}K^{-}$ $-\frac{\sqrt{3}}{2}F_{DP}$
@@
$T^{\prime+}_{c\bar{s}}\to D_{s}^{+}\eta$ $\frac{1}{\sqrt{2}}(c_{\phi^{\prime}}+\sqrt{2}s_{\phi^{\prime}})F_{DP}$ $T^{\prime+}_{c\bar{s}}\to D_{s}^{+}\eta^{\prime}$ $\frac{1}{\sqrt{2}}(s_{\phi^{\prime}}-\sqrt{2}c_{\phi^{\prime}})F_{DP}$
$16^{\prime}$ $T^{\prime\prime+}_{c}\to D^{0}\pi^{+}$ $\frac{1}{2}(F_{DP}+4AC_{P})$ $T^{\prime\prime+}_{c}\to D^{+}\pi^{0}$ $-\frac{1}{2\sqrt{2}}(F_{DP}+4AC_{P})$

### v3

@@ II.1 Scenarios I and II
[…]
Note that in this scenario, $T^{*}_{c\bar{s}0}(2900)^{++}$, $T_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ are in the $\overline{\bf 15}$. In particular, $T^{*}_{c\bar{s}0}(2900)^{++}$ and $T_{c\bar{s}0}(2900)^{0}$ are $T^{++}_{c\bar{s}}$ and $T^{0}_{c\bar{s}}$, respectively, belonging to the $S=1$ isotriplet, while $T^{*}_{cs0}(2870)^{0}$ is $T^{0}_{cs}$ belonging to another isotriplet with S=1S=-1. In this scenario, however, there are six other flavor exotic states similar to the above states, namely, $T^{+}_{cs}$. $T^{-}_{cs}$, $T^{++}_{c\bar{s}\bar{s}}$, $T^{+}_{c\bar{s}\bar{s}}$, $T^{++}_{c}$, and $T^{-}_{c}$. In fact, $T^{+}_{cs}$, $T^{-}_{cs}$ and $T^{*}_{cs0}(2870)^{0}$ belong to the same isotriplet (with S=1S=-1), while $T^{++}_{c\bar{s}\bar{s}}$, $T^{+}_{c\bar{s}\bar{s}}$, $T^{++}_{c}$ and $T^{-}_{c}$ belong to the S=2S=-2 isodouble and the iso-quarplet with $S=0$, respectively, as shown in Table 2. These nine flavor exotic states are labeled with asterisks in the table. They have been identified in the literature, see, for example, ref. Qin:2022nof , where a different notation was used. For convenience, we also show their notation for these states in Table 2.
Furthermore, as mentioned previously, mixing for particles in different isospin multiplets is expected to be more difficult. Consequently, $T^{+}_{c\bar{s}}$ in the $S=+1$ isotripet, $T^{+}_{c}$ and $T^{0}_{c}$ in the $S=0$ iso-quarplet do not mix with $T^{\prime\prime+}_{c\bar{s}}$, $D_{s}$, $T^{\prime+}_{c}$, $T^{\prime 0}_{c}$ or $D^{+,0}$. They should also be considered flavor exotic particles and be labeled with asterisks in the table. To our knowledge, these particles have not been considered flavor exotic particles in the literature. Altogether, we have twelve flavor exotic states in this scenario.
@@
where only the $\bar{\bf 3}^{\prime}$ states remain after taking the trace, $T^{i}_{\{ik\}}$, as the $\overline{\bf 15}$ is traceless, which can also be seen from Table 2.
@@ II.2 Remarks
In this work, we consider $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ either in the $\bf 6$} as in scenario I or in the $\overline{\bf 15}$ as in scenario II. In principle, they may belong to different multiplets. For example, we may have $T_{cs0}(2870)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{0}$ belonging to the $\bf 6$}, while $T_{cs0}(2870)^{0}$ belonging to the $\overline{\bf 15}$, or vice verse. It is our working assumption that they only belong to one of the multiplets, and we work out the consequences accordingly. It is useful to verify these predictions experimentally to check the working assumption. Therefore, we should pay attention to predictions where all $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ are involved, or, at least, $T_{cs0}(2870)^{0}$ and either one of $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ are involved.
As mentioned, it is our working assumption that particles in scenarios I and II do not mix. Even if they do, the amplitudes obtained in this work are still useful, since amplitudes for mixed states can be built upon them. Nevertheless, not all flavor exotic states can mix. Although $T^{++}_{c\bar{s}}$, $T^{0}_{c\bar{s}}$ and $T^{0}_{c\bar{s}}$ in scenarios I and II can, $T_{cs}^{0}$s cannot, as they have different isospin quantum numbers. In addition, the rest of the flavor exotic states in scenario II cannot, as they do not have any counterparts in scenario I, see Fig. 1. Furthermore, all flavor exotic states identified in this section do not mix with $D$ or any other $c\bar{q}$ states in any case. They are exotic.
@@ III.1 $T\to DP$ decays
We consider the strong decays of $T\to DP$ decays in this sub-section. We will first discuss how to decompose these decay amplitudes into topological amplitudes. Subsequently, $T\to DP$ decay amplitudes will be decomposed accordingly.
@@ III.1.1 Topological amplitudes in $T\to DP$ decays
[…]
$|F_{DP}|>|AC_{P}|,|AC_{D}|.$ (25)
The situation is similar to the strong decay of low-lying scalar particles, where a tetraquark configuration is favored Jaffe:1976ig ; Jaffe:1976ih , see also ref. Cheng:2002ai . For example, in the tetraquark picture, we have the following quark contents for $\sigma$ and $f_{0}$, $\sigma=ud\bar{u}\bar{d}$ and $f_{0}=(su\bar{s}\bar{u}+ds\bar{d}\bar{s})/\sqrt{2}$. The $\sigma\to\pi\pi$ decay amplitude is governed by a OZI superallowed fall-apart diagram, similar to $F_{DP}$ in Fig. 2 (a), while the $f_{0}\to\pi\pi$ decay amplitude is by a OZI allowed diagram, similar to $AC_{P}$ in Fig. 2 (b). Consequently, the $\sigma\to\pi\pi$ rate is larger than the $f_{0}\to\pi\pi$ rate, a fact supported by data ParticleDataGroup:2024cfk .
It is known that final state interaction can affect the size of a suppressed amplitude, see, for example, ref. Chua:2001br ; Chua:2005dt ; Chua:2007qw . Although $AC_{P}$ and $AC_{D}$ are expected to be smaller than $F_{DP}$, they can be enhanced to some extent through final state interaction. Their sizes remained to be determined experimentally. Hence, one should keep them in the $T\to DP$ decay amplitudes. On the other hand, from the above discussion, one may have an impression that in the presence of final state interaction, the above result, i.e., Eq. (25), based on the OZI rule, can be violated. This is, however, not true. For example, in $\overline{B}\to DP$ decays, one has exchange rescattering (governed by $r^{\prime}_{e}$) and annihilation rescattering (governed by $r^{\prime}_{a}$) in the presence of quasi-elastic rescattering Chua:2001br ; Chua:2005dt ; Chua:2007qw . The former is OZI allowed, while the latter is OZI suppressed. By fitting to data, it was found that $|r^{\prime}_{e}|>|r^{\prime}_{a}|$ Chua:2005dt ; Chua:2007qw , hence, in line with the OZI rule. Therefore, final state interaction will not alter Eq. (25), as they both respect the OZI rule.
The decay amplitudes of tetraquarks in the $\bf 6$} in scenario I and in the $\overline{\bf 15}$ in scenario II can only have contributions from diagram (a), the fall-apart amplitude, $F_{DP}$. This can be easily seen, as the $AC_{P}$ and $AC_{D}$ terms in Eqs. (16) and (17) necessary involve the trace $T^{i}_{[ik]}$ or $T^{i}_{\{ik\}}$, where only the states in the $\bar{\bf 3}$ in scenario I and the $\bar{\bf 3}^{\prime}$ in scenario II remain, see Eqs.(11) and (15). In fact, the reason for the above facts can be easily understood by noting $\bar{\bf 3}\otimes{\bf 8}=\overline{\bf 15}\oplus{\bf 6}\oplus\bar{\bf 3}$, see Eq. (176). Hence there is only one $\overline{\bf 15}$ and one $\bf 6$} in $D(\bar{\bf 3})\Pi({\bf 8})$ final states, which can match the $\overline{\bf 15}$ of $T$ in scenario II and $\bf 6$} of $T$ in scenario I, respectively. Adding $\eta_{1}$ in $P$ does not affect the above matching.
@@ III.1.2 $T\to DP$ decay amplitudes
$1^{*}$ $T^{++}_{c\bar{s}}\to D^{+}K^{+}$ $F_{DP}$ $T^{++}_{c\bar{s}}\to D_{s}^{+}\pi^{+}$ $F_{DP}$
$2^{*}$ $T^{+}_{c\bar{s}}\to D^{0}K^{+}$ $\frac{1}{\sqrt{2}}F_{DP}$ $T^{+}_{c\bar{s}}\to D^{+}K^{0}$ $\frac{1}{\sqrt{2}}F_{DP}$
@@
$3^{*}$ $T^{0}_{c\bar{s}}\to D^{0}K^{0}$ $F_{DP}$ $T^{0}_{c\bar{s}}\to D_{s}^{+}\pi^{-}$ $F_{DP}$
$4^{*}$ $T^{0}_{cs}\to D^{0}\overline{K}^{0}$ $F_{DP}$ $T^{0}_{cs}\to D^{+}K^{-}$ $F_{DP}$
5 $T^{+}_{c}\to D^{0}\pi^{+}$ $\frac{1}{\sqrt{2}}F_{DP}$ $T^{+}_{c}\to D^{+}\pi^{0}$ $\frac{1}{2}F_{DP}$
$T^{+}_{c}\to D_{s}^{+}\overline{K}^{0}$ $\frac{1}{\sqrt{2}}F_{DP}$ $T^{+}_{c}\to D^{+}\eta$ $\frac{c_{\theta}+\sqrt{2}s_{\theta}}{2}F_{DP}$
$T^{+}_{c}\to D^{+}\eta^{\prime}$ $\frac{\sqrt{2}c_{\theta}-s_{\theta}}{2}F_{DP}$
6 $T^{0}_{c}\to D^{0}\pi^{0}$ $\frac{1}{2}F_{DP}$ $T^{0}_{c}\to D^{+}\pi^{-}$ $\frac{F_{DP}}{\sqrt{2}}$
$T^{0}_{c}\to D_{s}^{+}K^{-}$ $\frac{1}{\sqrt{2}}F_{DP}$ $T^{0}_{c}\to D^{+}\eta$ $\frac{c_{\theta}+\sqrt{2}s_{\theta}}{2}F_{DP}$
$T^{0}_{c}\to D^{+}\eta^{\prime}$ $\frac{\sqrt{2}c_{\theta}-s_{\theta}}{2}F_{DP}$
7 $T^{\prime\prime+}_{c}\to D^{0}\pi^{+}$ $\frac{1}{\sqrt{2}}(F_{DP}+2AC_{P})$ $T^{\prime\prime+}_{c}\to D^{+}\pi^{0}$ $\frac{1}{2}(F_{DP}+2AC_{P})$
@@ 6 $T^{0}_{c}\to D^{0}\pi^{0}$ $\frac{1}{2}F_{DP}$ $T^{0}_{c}\to D^{+}\pi^{-}$ $\frac{F_{DP}}{\sqrt{2}}$
$T^{\prime\prime+}_{c}\to D^{+}\eta$ $\frac{c_{\theta}+\sqrt{2}s_{\theta}}{6}(F_{DP}+2AC_{P})$ $T^{\prime\prime+}_{c}\to D^{+}\eta^{\prime}$ $\frac{s_{\theta}-\sqrt{2}c_{\theta}}{6}(F_{DP}+2AC_{P})$
$\frac{\sqrt{2}s_{\theta}-2c_{\theta}}{3}(F_{DP}-AC_{P}-3AC_{D})$ $\frac{\sqrt{2}c_{\theta}+2s_{\theta}}{3}(F_{DP}-AC_{P}-3AC_{D})$
@@
$T^{\prime\prime 0}_{c}\to D^{0}\eta$ $\frac{c_{\theta}+\sqrt{2}s_{\theta}}{6}(F_{DP}+2AC_{P})$ $T^{\prime\prime 0}_{c}\to D^{+}\eta^{\prime}$ $\frac{s_{\theta}-\sqrt{2}c_{\theta}}{6}(F_{DP}+2AC_{P})$
$\frac{\sqrt{2}s_{\theta}-2c_{\theta}}{3}(F_{DP}-AC_{P}-3AC_{D})$ $\frac{\sqrt{2}c_{\theta}+2s_{\theta}}{3}(F_{DP}-AC_{P}-3AC_{D})$
9 $T^{\prime\prime+}_{c\bar{s}}\to D^{0}K^{+}$ $\frac{1}{\sqrt{2}}(F_{DP}+2AC_{P})$ $T^{\prime\prime+}_{c\bar{s}}\to D^{+}K^{0}$ $\frac{1}{\sqrt{2}}(F_{DP}+2AC_{P})$
$T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\eta$ $\frac{c_{\theta}+\sqrt{2}s_{\theta}}{3}(F_{DP}+2AC_{P})$ $T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\eta^{\prime}$ $\frac{\sqrt{2}c_{\theta}-s_{\theta}}{3}(F_{DP}+2AC_{P})$
$\frac{\sqrt{2}s_{\theta}-2c_{\theta}}{3}(F_{DP}-AC_{P}-3AC_{D})$ $\frac{\sqrt{2}c_{\theta}+2s_{\theta}}{3}(F_{DP}-AC_{P}-3AC_{D})$
@@
$2^{\prime*}$ $T^{+}_{c\bar{s}}\to D^{0}K^{+}$ $\frac{1}{\sqrt{2}}F_{DP}$ $T^{+}_{c\bar{s}}\to D^{+}K^{0}$ $\frac{1}{\sqrt{2}}F_{DP}$
@@
$11^{\prime*}$ $T^{0}_{c}\to D^{0}\pi^{0}$ $\frac{2}{\sqrt{3}}F_{DP}$ $T^{0}_{c}\to D^{+}\pi^{-}$ $\sqrt{\frac{2}{3}}F_{DP}$
@@
$13^{\prime}$ $T^{\prime+}_{c}\to D^{0}\pi^{+}$ $\frac{1}{2\sqrt{3}}F_{DP}$ $T^{\prime+}_{c}\to D^{+}\pi^{0}$ $\frac{1}{2\sqrt{6}}F_{DP}$
$T^{\prime+}_{c}\to D^{+}\eta$ $\frac{\sqrt{6}}{4}(c_{\theta}+\sqrt{2}s_{\theta})F_{DP}$ $T^{\prime+}_{c}\to D^{+}\eta^{\prime}$ $\frac{\sqrt{6}}{4}(s_{\theta}-\sqrt{2}c_{\theta})F_{DP}$
$T^{\prime+}_{c}\to D_{s}^{+}\overline{K}^{0}$ $\frac{\sqrt{3}}{2}F_{DP}$
$14^{\prime}$ $T^{\prime 0}_{c}\to D^{0}\pi^{0}$ $\frac{1}{2\sqrt{6}}F_{DP}$ $T^{\prime 0}_{c}\to D^{0}\eta$ $\frac{\sqrt{6}}{4}(c_{\theta}+\sqrt{2}s_{\theta})F_{DP}$
$T^{\prime 0}_{c}\to D^{0}\eta^{\prime}$ $\frac{\sqrt{6}}{4}(s_{\theta}-\sqrt{2}c_{\theta})F_{DP}$ $T^{\prime 0}_{c}\to D^{+}\pi^{-}$ $\frac{1}{2\sqrt{3}}F_{DP}$
$T^{\prime 0}_{c}\to D_{s}^{+}K^{-}$ $\frac{\sqrt{3}}{2}F_{DP}$
@@
$T^{\prime+}_{c\bar{s}}\to D_{s}^{+}\eta$ $\frac{1}{\sqrt{2}}(c_{\theta}+\sqrt{2}s_{\theta})F_{DP}$ $T^{\prime+}_{c\bar{s}}\to D_{s}^{+}\eta^{\prime}$ $\frac{1}{\sqrt{2}}(s_{\theta}-\sqrt{2}c_{\theta})F_{DP}$
$16^{\prime}$ $T^{\prime\prime+}_{c}\to D^{0}\pi^{+}$ $\frac{1}{2}(F_{DP}+4AC_{P})$ $T^{\prime\prime+}_{c}\to D^{+}\pi^{0}$ $\frac{1}{2\sqrt{2}}(F_{DP}+4AC_{P})$